Johnson Tire Distributors has an unlevered cost of capital of 10 percent, a tax rate of 34 percent, and expected earnings before interest and taxes of $1,700. The company has $3,200 in bonds outstanding that have an 8 percent coupon and pay interest annually. The bonds are selling at par value. What is the cost of equity? 8.37 percent 7.32 percent 9.42 percent 10.46 percent 6.28 percent
Answer #1

D. 10.46 percent

VU = [$1,700 x (1 - 0.34)] / 0.10 = $11,220

VL = $11,220 + (0.34 x $3,200) = $12,308

VE = $12,308 - $3,200 = $9,108

RE = 0.10 + [(0.10 - 0.08) x ($3,200/$9,108) x (1 - 0.34)] = 10.46 percent